At the city museum, child admission is $5.30 and adult admission is $9.00. On Saturday, 136 tickets were sold for a total sales of $913.20. How many child tickets were sold that day?
Found 2 solutions by richwmiller, ikleyn:
Answer by richwmiller(17219)

c+a=136
5.3*c+9*(a)=913.20
5.3*c+9*(136-c)=913.20

Answer by ikleyn(52781)

.
Your equations are

     c +      a = 136,       (1)
5.30*c + 9.00*a = 913.20.    (2)

From equation (1), express a = 136 - c and substitute it into equation (2) replacing  "a". You will get a single equation for "c":

5.30*c + 9*(136-c) = 913.20.


Simplify and solve for "c", the number of children's tickets.

This method is called "the Substitution method".